OpenStudy (agentjamesbond007):

If f ' (x) = |x-2|, what is a possible graph of y = f(x)

OpenStudy (agentjamesbond007):

please draw

OpenStudy (blockcolder):

\[\large f'(x)= \begin{cases} x-2 &\text{if }x\geq2\\ 2-x &\text{if }x<2 \end{cases}\\ \large f(x)= \begin{cases} \frac{1}{2}x^2-2x+c_1 &\text{if }x\geq 2\\ -\frac{1}{2}x^2+2x+c_2 &\text{if }x<2 \end{cases}\] where c1 and c2 are arbitrary constants. You can draw these, right?

OpenStudy (agentjamesbond007):

So they would have different limits at x=2?

OpenStudy (blockcolder):

That depends on the constants. What you can do is just pick convenient values for c1 and c2 and draw the graphs accordingly.

OpenStudy (anonymous):

Well, since f' is an absolute value equation, it will always be non-negative, and you can see that the slope is 0 when x=2.

OpenStudy (agentjamesbond007):

 Here are my options